OpenStudy (anonymous):

(1-(2/3x))/(x-(4/9x)) simplify steps and explanation please

OpenStudy (anonymous):

\[\frac{1-\frac{2}{3x}}{x-\frac{4}{9x}}\] multiply top and bottom by 9x to clear the fraction in one step

OpenStudy (anonymous):

\[\frac{1-\frac{2}{3x}}{x-\frac{4}{9x}}\times \frac{9x}{9x}\] \[=\frac{9x-6}{9x^2-4}\]

OpenStudy (anonymous):

you can factor both top and bottom but nothing cancels

OpenStudy (anonymous):

oh maybe it does

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

\[\frac{3(2x-2)}{(3x+2)(3x-2)}\] nope nothing cancels yw